# Fractional diffusion as the limit of a short range potential Rayleigh gas

### v1

@@ 4.6. The Empirical Distribution
$=P_{t}(S\setminus\mathcal{G(\varepsilon)})+\underbrace{P_{t}(S\cap\mathcal{G(\varepsilon)})-\hat{P_{t}}(S\cap\mathcal{G(\varepsilon)})}_{\leq 0}$
$\leq P_{t}(\mathcal{MT}\setminus\mathcal{G(\varepsilon)}).$ (4.30)
Notice that $\varepsilon$ does not depend on $S$, so this holds true for every $S\subset\mathcal{MT}$ and for any $S\subset\mathcal{MT}$ we also have as in (4.30)
@@
$\leq P_{t}(\mathcal{MT}\setminus\mathcal{G(\varepsilon)})$ (4.31)
@@
$\quad+C(M(\varepsilon))^{2}\left(\frac{\tilde{\eta}V(\varepsilon)\varepsilon}{\tilde{\delta}^{2}}+\tilde{\delta}+\frac{TV(\varepsilon)}{\tilde{\eta}}\left(\frac{\varepsilon}{\tilde{\eta}}\right)^{2}\right)$
$\quad+\frac{M_{f_{0}}}{V(\varepsilon)}+(C_{g}+\tilde{C}_{q,n}ct)\frac{M(\varepsilon)}{V(\varepsilon)}.$ (4.32)
@@ 4.7. Making time $T$ large
The aim of this section is to make time $T$ large, as $\varepsilon$ tends to be very small. We are doing this by equating the leading order terms of the bound of the probability of bad trees in (4.19) and making them small. Therefore we have the following lemma.
@@
$c^{\frac{28}{29}}T=\varepsilon^{\frac{38m}{29}-\frac{5}{29}}$
which becomes large, as $\varepsilon$ becomes small, for the values of $m$ such that $0<m<\frac{5}{38}.$
@@
In this step, we want to make the time $cT$ large such that $\lim_{\varepsilon\to 0}cT(\varepsilon)=\infty.$ By considering only the leading order terms in (4.19) we have
$\frac{cT}{M(\varepsilon)}=M^{2}(\varepsilon)\tilde{\delta}=\frac{M^{2}(\varepsilon)TV(\varepsilon)\varepsilon^{2}}{\tilde{\eta}^{3}}=\frac{\varepsilon M^{2}(\varepsilon)V(\varepsilon)\tilde{\eta}}{\tilde{\delta}^{2}}=\frac{cTM(\varepsilon)}{V(\varepsilon)}.$ (4.33)
@@
then consider
$\frac{\varepsilon M^{2}(\varepsilon)V(\varepsilon)\tilde{\eta}}{\tilde{\delta}^{2}}=\frac{M^{2}(\varepsilon)TV(\varepsilon)\varepsilon^{2}}{\tilde{\eta}^{3}}\Leftrightarrow\tilde{\eta}=T^{\frac{1}{4}}\varepsilon^{\frac{1}{4}}\tilde{\delta}^{\frac{1}{2}}$ (4.35)
and
$M^{2}(\varepsilon)\tilde{\delta}=\frac{\varepsilon M^{2}(\varepsilon)V(\varepsilon)\tilde{\eta}}{\tilde{\delta}^{2}}$ $\Leftrightarrow\tilde{\delta}^{3}=\varepsilon V(\varepsilon)\tilde{\eta}=\varepsilon V(\varepsilon)T^{\frac{1}{4}}\varepsilon^{\frac{1}{4}}\tilde{\delta}^{\frac{1}{2}}$
$\Leftrightarrow\tilde{\delta}=\varepsilon^{\frac{1}{2}}(V(\varepsilon))^{\frac{2}{5}}T^{\frac{1}{10}}.$ (4.36)
Then with (4.34) we have
$\frac{cT}{M(\varepsilon)}=M^{2}(\varepsilon)\tilde{\delta}$ $\Leftrightarrow cT=M^{3}(\varepsilon)\varepsilon^{\frac{1}{2}}(V(\varepsilon))^{\frac{2}{5}}T^{\frac{1}{10}}$
$\Leftrightarrow V(\varepsilon)^{\frac{19}{10}}=c\varepsilon^{-\frac{1}{2}}T^{\frac{9}{10}}$
$\Leftrightarrow V(\varepsilon)=c^{\frac{10}{19}}\varepsilon^{-\frac{5}{19}}T^{\frac{9}{19}}.$ (4.37)
@@
$\frac{cT}{M(\varepsilon)}=\frac{cT}{\sqrt{V(\varepsilon)}}=c^{\frac{28}{38}}T^{\frac{29}{38}}\varepsilon^{\frac{5}{38}}.$
@@
$\varepsilon^{m}=c^{\frac{28}{38}}T^{\frac{29}{38}}\varepsilon^{\frac{5}{38}}\Leftrightarrow c^{\frac{28}{29}}T=\varepsilon^{\frac{38m}{29}-\frac{5}{29}}.$ (4.38)
Next, we find the values of $m$ so that $c^{\frac{28}{29}}T$ is large. Thus it needs to satisfy
$\frac{38m}{29}<\frac{5}{29}\Leftrightarrow 0<m<\frac{5}{38}.$
Then it can be easily checked that all further terms in (4.29) are of higher order by choosing $\delta=\sqrt{\varepsilon}$. ∎
@@ 5. From Linear Boltzmann to the Fractional diffusion equation
This section is devoted in the proof of Theorem 1.2. We are doing so by adapting results from [MMM11] and combine them with Theorem 1.1. First, we are splitting the linear Boltzmann operator into the gain and loss term and write the gain term via the Carleman representation in order to integrate over the $v_{2}^{\prime}\in E_{v_{1},v_{1}^{\prime}}$ and $v_{1}^{\prime}\in\mathds{R}^{3}$ variables instead of $\omega\in\mathbb{S}^{2}$ and $v_{2}\in\mathds{R}^{3}$. Then, we are finding the behaviour of the gain term and loss term, $K(f)(v_{1})$ and $\nu(v_{1})$ respectively, for large velocities $|v_{1}|$ and for the case $\Theta$ close to zero. Additionally, we make the right choices of $\alpha$ and $\beta$ in order to apply [MMM11, Theorem 3.2], which is a theorem which derives a Fractional diffusion equation from the Linear Boltzmann equation. Finally, we combine this result with Theorem 1.1 and obtain the proof of the full derivation, i.e. the proof of Theorem 1.2.
The gain term of the linear Boltzmann collision operator is
$Q^{+}(f)(v_{1})=\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}f(v_{1}^{\prime})g_{0}(v_{2}^{\prime})\ B(v_{1}-v_{2},\omega)\,\mathrm{d}\omega\,\mathrm{d}v_{2},$
@@
which represents the probability that a particle after collision has velocity $v_{1}$. Furthermore, $\omega$ is the orthonormal vector and gives the direction of the apse line, which is the line that bisects the angle between $v_{1}-v_{2}$ and $v_{1}^{\prime}-v_{2}^{\prime}$. The idea of Carleman representation, first introduced by Carleman [Car57], is to parameterise the gain term $Q^{+}$ by the variables $v_{1}^{\prime}$, $v_{2}^{\prime}$ instead of $v_{2}$, $\omega$. Rather than parameterising the sphere $\mathbb{S}_{v_{1},v_{2}}$, one considers the set $E_{v_{1},v_{1}^{\prime}}$ which is the set of all $v_{2}^{\prime}$ which could be the velocity after collision of one particle, if the other velocity after collision is $v_{1}^{\prime}$ and one of the velocities before collision is $v_{1}$. The set $E_{v_{1},v_{1}^{\prime}}$ is the hyperplane orthogonal to $v_{1}-v_{1}^{\prime}$ and containing $v_{1}$ and we are writing it as the set $E_{v_{1},v_{1}^{\prime}}:=\{v_{2}^{\prime}\in\mathds{R}^{3}\ :\ (v_{2}^{\prime}-v_{1})\cdot(v_{1}-v_{1}^{\prime})=0\}$, where $(v_{2}-v_{1})\cdot(v_{1}-v_{1}^{\prime})$ is the usual inner product $(v_{2}-v_{1})\cdot(v_{1}-v_{1}^{\prime})=\langle v_{2}-v_{1},v_{1}-v_{1}^{\prime}\rangle=|v_{2}-v_{1}||v_{1}-v_{1}^{\prime}|\cos(v_{2}-v_{1},v_{1}-v_{1}^{\prime})$. See Pulvirenti, Wennberg [PW97, Section 3] for a further discussion of this. In order to represent the gain term with the Carleman representation, we set $v_{1}^{\prime}=v_{1}+q\omega$, $q\in\mathds{R}$ and $\omega\in\mathbb{S}^{2}$. Now, for any fixed $\omega$ and $v_{1}$ we can write $v_{2}^{\prime}=v_{2}-q\omega$, so $v_{2}=v_{2}^{\prime}+q\omega$. Then, with this change of variables, $\mathrm{d}v_{2}=\mathrm{d}v_{2}^{\prime}\mathrm{d}q$. Furthermore, by taking the polar coordinates of $v_{1}^{\prime}$ centred at $v_{1}$ we get
@@
$Q^{+}(f)(v_{1})=2\int_{\mathds{R}^{3}}\frac{f(v_{1}^{\prime})}{|v_{1}-v_{1}^{\prime}|^{2}}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})\ B(v_{1}-v_{2},\omega)\,\mathrm{d}v_{2}^{\prime}\,\mathrm{d}v_{1}^{\prime}.$
Now in order to have the variables $v_{1}^{\prime}$, $v_{2}^{\prime}$ inside the collision kernel $B$, we use the identity $v_{1}-v_{2}=2v_{1}-v_{1}^{\prime}-v_{2}^{\prime}$ which comes from the conservation of momentum and we take
$Q^{+}(f)(v_{1})=2\int_{\mathds{R}^{3}}\frac{f(v_{1}^{\prime})}{|v_{1}-v_{1}^{\prime}|^{2}}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})\ B(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}\,\mathrm{d}v_{1}^{\prime}.$
@@
$Q(f)(v_{1})$ $=\int_{\mathds{R}^{3}}[\sigma(v_{1},v_{1}^{\prime})f(v_{1}^{\prime})-\sigma(v_{1}^{\prime},v_{1})f(v_{1})]\,\mathrm{d}v_{1}^{\prime}=K(f)(v_{1})-\nu(v_{1})f(v_{1}),$ (5.1)
@@
$=\int_{\mathds{R}^{3}}\left[\frac{2}{|v_{1}-v_{1}^{\prime}|^{2}}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})B(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}\right]f(v_{1}^{\prime})\,\mathrm{d}v_{1}^{\prime}$
@@
5.1. Estimate the gain term of the collision operator
By the above, the Carleman representation of the gain term of the linear Boltzmann operator $Q$ is given by
$K(f)(v_{1})$ $=\int_{\mathds{R}^{3}}\left[\frac{2}{|v_{1}-v_{1}^{\prime}|^{2}}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})B(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}\right]f(v_{1}^{\prime})\,\mathrm{d}v_{1}^{\prime},$
where $E_{v_{1},v_{1}^{\prime}}$ is the set of all $v_{2}^{\prime}$ which could be the velocity after collision of one particle, if the other velocity after collision is $v_{1}^{\prime}$ and one of the velocities before collision is $v_{1}$. The set $E_{v_{1},v_{1}^{\prime}}$ is the hyperplane orthogonal to $v_{1}-v_{1}^{\prime}$ which contains $v_{1}$ and it is defined as the set $E_{v_{1},v_{1}^{\prime}}:=\{v_{2}^{\prime}\in\mathds{R}^{3}\ :\ (v_{2}^{\prime}-v_{1})\cdot(v_{1}-v_{1}^{\prime})=0\}$. Let us now estimate the integrand of $K(f)(v_{1})$ when $\Theta\in(0,\frac{\pi}{2}-\sigma)$. That is
$\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})B$ $(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\bigg{[}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})|2v_{1}-v_{1}^{\prime}-v_{2}^{\prime}|^{\frac{n-5}{n-1}}\Theta(\sin\Theta)^{-1}\,\mathrm{d}v_{2}^{\prime}$
$\quad+\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})\mathcal{O}(|2v_{1}-v_{1}^{\prime}-v_{2}^{\prime}|^{\frac{n-7}{n-1}})\Theta(\sin\Theta)^{-1}\,\mathrm{d}v_{2}^{\prime}\bigg{]}.$
We know that
$\Leftrightarrow|v_{1}-v_{2}^{\prime}|^{2}=\tan^{2}\Theta|v_{1}-v_{1}^{\prime}|^{2}$
where we used the fact that $(v_{1}-v_{2}^{\prime})\cdot(v_{1}-v_{1}^{\prime})=0$, since $v_{2}^{\prime}\in E_{v_{1},v_{1}^{\prime}}$. Furthermore, the angle $\Theta$ is bounded by $\frac{\pi}{2}-\sigma$ and the function $\tan x$ is increasing in $(0,\frac{\pi}{2}-\sigma)$, thus we obtain
$|v_{1}-v_{2}^{\prime}|\leq\tan(\frac{\pi}{2}-\sigma)|v_{1}-v_{1}^{\prime}|,$
i.e. $v_{2}^{\prime}\in B(v_{1},\tan(\frac{\pi}{2}-\sigma)|v_{1}-v_{1}^{\prime}|).$ Hence, by using polar coordinates, the integral becomes
$\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})B(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\bigg{[}\int_{0}^{\tan(\frac{\pi}{2}-\sigma)|v_{1}-v_{1}^{\prime}|}\int_{0}^{2\pi}g_{0}(v_{1}+rv_{s})|2v_{1}-v_{1}^{\prime}-v_{2}^{\prime}|^{\frac{n-5}{n-1}}\Theta(r)(\sin\Theta(r))^{-1}\,\mathrm{d}s\,r\,\mathrm{d}r$
$\quad+\int_{0}^{\tan(\frac{\pi}{2}-\sigma)|v_{1}-v_{1}^{\prime}|}\int_{0}^{2\pi}g_{0}(v_{1}+rv_{s})\mathcal{O}(|2v_{1}-v_{1}^{\prime}-v_{2}^{\prime}|^{\frac{n-7}{n-1}})\Theta(r)(\sin\Theta(r))^{-1}\,\mathrm{d}s\,r\,\mathrm{d}r\bigg{]}$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\bigg{[}\int_{0}^{\tan(\frac{\pi}{2}-\sigma)|v_{1}-v_{1}^{\prime}|}\int_{0}^{2\pi}g_{0}(v_{1}+rv_{s})(|v_{1}-v_{1}^{\prime}|^{2}+r^{2})^{\frac{n-5}{2(n-1)}}\Theta(r)(\sin\Theta(r))^{-1}\,\mathrm{d}s\,r\,\mathrm{d}r$
$\quad+\int_{0}^{\tan(\frac{\pi}{2}-\sigma)|v_{1}-v_{1}^{\prime}|}\int_{0}^{2\pi}g_{0}(v_{1}+rv_{s})\mathcal{O}((|v_{1}-v_{1}^{\prime}|^{2}+r^{2})^{\frac{n-7}{2(n-1)}})\Theta(r)(\sin\Theta(r))^{-1}\,\mathrm{d}s\,r\,\mathrm{d}r\bigg{]}.$
Now, we use the change of variables $r:=|v_{1}-v_{1}^{\prime}|\rho$ and we obtain
$\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}|v_{1}-v_{1}^{\prime}|^{2}\bigg{[}\int_{0}^{\tan(\frac{\pi}{2}-\sigma)}\big{[}|v_{1}$ $-v_{1}|^{\frac{n-5}{n-1}}\rho(1+\rho^{2})^{\frac{n-5}{2(n-1)}}+\mathcal{O}(|v_{1}-v_{1}|^{\frac{n-7}{n-1}})\rho(1+\rho^{2})^{\frac{n-7}{2(n-1)}}\big{]}$
×$\times\frac{\Theta(\rho|v_{1}-v_{1}^{\prime}|)}{\sin\Theta(\rho|v_{1}-v_{1}^{\prime}|)}\int_{0}^{2\pi}g_{0}(v_{1}+|v_{1}-v_{1}^{\prime}|\rho v_{s})\,\mathrm{d}s\,\mathrm{d}\rho\bigg{]}.$
Now, we observe that
which implies that
$\tan\Theta=\rho\Leftrightarrow\Theta=\arctan\rho.$
where
$\left|\frac{v_{1}}{|v_{1}|}+\frac{|v_{1}-v_{1}^{\prime}|}{|v_{1}|}\rho v_{s}\right|^{2}=1+\frac{|v_{1}-v_{1}^{\prime}|^{2}}{|v_{1}|^{2}}\rho^{2}+2\rho\frac{|v_{1}-v_{1}^{\prime}|}{|v_{1}|}\frac{v_{1}}{|v_{1}|}\cdot v_{s},$
which converges to $1+\rho^{2}$, as $|v_{1}|\to\infty$. Here we used the fact that $v_{s}\cdot(v_{1}-v_{1}^{\prime})=0.$ Therefore,
$\int_{0}^{2\pi}g_{0}(v_{1}+|v_{1}-v_{1}^{\prime}|\rho v_{s})\,\mathrm{d}s=2\pi C_{2}|v_{1}|^{-q}\left((1+\rho^{2})^{-\frac{q}{2}}+\mathcal{O}(\frac{1}{|v_{1}|})\right),\quad as\ |v_{1}|\to\infty.$
Hence, the whole integral becomes
$\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})B(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}2\pi C_{2}|v_{1}|^{-q}|v_{1}-v_{1}^{\prime}|^{2}$
×$\quad\times\bigg{[}|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}\int_{0}^{\tan(\frac{\pi}{2}-\sigma)}\rho(1+\rho^{2})^{\frac{n-5}{2(n-1)}-\frac{q}{2}}\frac{\arctan\rho}{\sin(\arctan\rho)}+\mathcal{O}(\frac{1}{|v_{1}|})\mathrm{d}\rho$
$\qquad+\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\int_{0}^{\tan(\frac{\pi}{2}-\sigma)}\rho(1+\rho^{2})^{\frac{n-7}{2(n-1)}-\frac{q}{2}}\frac{\arctan\rho}{\sin(\arctan\rho)}+\mathcal{O}(\frac{1}{|v_{1}|})\mathrm{d}\rho\bigg{]}$
$=:\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}2\pi C_{2}|v_{1}|^{-q}|v_{1}-v_{1}^{\prime}|^{2}\bigg{[}|v_{1}-v_{1}^{\prime}|^{\frac{n-5}{n-1}}C_{S}+\mathcal{O}(|v_{1}-v_{1}^{\prime}|^{\frac{n-7}{n-1}})\bigg{]},$
as $|v_{1}|\to\infty$. We state now the following proposition which will be used later for finding the behaviour of the gain term for large velocities. We observe that when $n=5$, there is nothing to prove.
Proposition 5.1.

### v2

@@ 4.8. The Empirical Distribution
$=P_{t}(S\setminus\mathcal{G(\varepsilon)})+\underbrace{P_{t}(S\cap\mathcal{G(\varepsilon)})-\hat{R_{t}}(S\cap\mathcal{G(\varepsilon)})}_{\leq 0}+(\hat{R}_{t}(S\cap\mathcal{G(\varepsilon)})-\hat{P_{t}}(S\cap\mathcal{G(\varepsilon)}))$
$\leq P_{t}(\mathcal{MT}\setminus\mathcal{G(\varepsilon)})+(1-\zeta(\varepsilon)).$ (4.33)
We obtain further bounds by applying (4.33) to complements.
@@
$\leq P_{t}(\mathcal{MT}\setminus\mathcal{G(\varepsilon)})+(1-\zeta(\varepsilon)).$ (4.34)
@@
[…]
$\quad+\frac{M_{f_{0}}}{V(\varepsilon)}+(C_{g}+\tilde{C}_{q,n}ct)\frac{M(\varepsilon)}{V(\varepsilon)}+CM(\varepsilon)(I(\varepsilon)+(J(\varepsilon)^{3}).$ (4.35)
@@ 4.9. Making time $T$ large
The aim of this section is to make time $T$ large, as $\varepsilon$ tends to be very small. We are doing this by equating the leading order terms of the bound of the probability of bad trees in (4.32) and making them small whereas also ensuring that (4.28) holds. Therefore we have the following lemma.
@@
$cT=\varepsilon^{\frac{4m}{3}-\frac{2}{9}}$
which becomes large, as $\varepsilon$ becomes small, for the values of $m$ such that $0<m<\frac{1}{6}.$
@@
In this step, we want to make the time $cT$ large such that $\lim_{\varepsilon\to 0}cT(\varepsilon)=\infty.$ By considering only the leading order terms in (4.32) and for $\tilde{}\eta=\tilde{\delta}$ we have
$\frac{cT}{M(\varepsilon)}=M(\varepsilon)\tilde{\delta}=\frac{M^{2}(\varepsilon)V(\varepsilon)\varepsilon}{\sqrt{\tilde{\delta}}}=\frac{cTM(\varepsilon)}{V(\varepsilon)}.$ (4.36)
@@
and consider
$M(\varepsilon)\tilde{\delta}=\frac{\varepsilon M^{2}(\varepsilon)V(\varepsilon)}{\sqrt{\tilde{\delta}}}$ $\Leftrightarrow\tilde{\delta}^{\frac{3}{2}}=\varepsilon M^{3}(\varepsilon)\Leftrightarrow\tilde{\delta}=\varepsilon^{\frac{2}{3}}M^{2}(\varepsilon).$ (4.38)
Then with (4.37) and (4.38) we have
$\frac{cT}{M(\varepsilon)}=M(\varepsilon)\tilde{\delta}$ $\Leftrightarrow cT=M^{4}(\varepsilon)\varepsilon^{\frac{2}{3}}$
$\Leftrightarrow M(\varepsilon)=(cT)^{\frac{1}{4}}\varepsilon^{-\frac{1}{6}}.$ (4.39)
@@
$\frac{cT}{M(\varepsilon)}=(cT)^{\frac{3}{4}}\varepsilon^{\frac{1}{6}}.$
@@
$\varepsilon^{m}=(cT)^{\frac{3}{4}}\varepsilon^{\frac{1}{6}}\Leftrightarrow cT=\varepsilon^{\frac{4m}{3}-\frac{2}{9}}.$ (4.40)
Next, we find the values of $m$ so that $cT$ is large. Thus it needs to satisfy
$\frac{4m}{3}<\frac{2}{9}\Leftrightarrow 0<m<\frac{1}{6}.$
The term $0<1-\zeta(\varepsilon)\leq Cc\varepsilon$ is smaller than $\varepsilon^{1/6}$ for any admissible choice of $c$. It can be easily checked that all further terms in (4.32) are of higher order by choosing $\delta=\sqrt{\varepsilon}$. Similarly, if $I(\varepsilon)=\sqrt{\varepsilon}$ and $J(\varepsilon)=\varepsilon^{1/3}$ then the $\textnormal{Tri}(\varepsilon)$ and $\textnormal{Slow}(\varepsilon)$ are also of higher order compared to the $\textnormal{Rec}(\varepsilon)$. We note that with this choice (4.28) holds. ∎
@@ 5. From Linear Boltzmann to the fractional diffusion equation
The aim of this section is to give the proof of Theorem 1.2. We are doing so by adapting results from [MMM11] and combine them with Theorem 1.1. First, we are splitting the linear Boltzmann operator into the gain and loss term and write the gain term via the Carleman representation in order to integrate over the $v_{2}^{\prime}\in E_{v_{1},v_{1}^{\prime}}$ and $v_{1}^{\prime}\in\mathds{R}^{3}$ variables instead of $\omega\in\mathbb{S}^{2}$ and $v_{2}\in\mathds{R}^{3}$. Then, we are finding the behaviour of the gain and loss term, $K(f)(v_{1})$ and $\nu(v_{1})$ respectively, for large velocities $|v_{1}|$ and for $\Theta\in(0,\frac{\pi}{2})$. Additionally, we appropriately choose $\alpha$ and $\beta$ so that [MMM11, Theorem 3.2] can be applied, which is a theorem which derives a fractional diffusion equation from the linear Boltzmann equation. Finally, we combine this result with Theorem 1.1 and obtain the proof of the full derivation, i.e. the proof of Theorem 1.2. The gain term of the linear Boltzmann collision operator is
$Q^{+}(f)(v_{1})=\int_{\mathds{R}^{3}}\int_{\mathbb{S}^{2}}f(v_{1}^{\prime})g_{0}(v_{2}^{\prime})\ \mathcal{B}(v_{1}-v_{2},\omega)\,\mathrm{d}\omega\,\mathrm{d}v_{2},$
@@
which represents the probability that a particle after collision has velocity $v_{1}$. Furthermore, $\omega$ is the orthonormal vector and gives the direction of the apse line, which is the line that bisects the angle between $v_{1}-v_{2}$ and $v_{1}^{\prime}-v_{2}^{\prime}$. The idea of Carleman representation, first introduced by Carleman [Car57], is to parameterise the gain term $Q^{+}$ by the variables $v_{1}^{\prime}$, $v_{2}^{\prime}$ instead of $v_{2}$, $\omega$. Rather than parameterising the sphere $\mathbb{S}_{v_{1},v_{2}}$, which is defined as the set of all possible $v^{\prime}$ which result from a collision between one particle with velocity $v_{1}$ and another particle with velocity $v_{2}$, i.e. the sphere with a diagonal the line segment which connects $v_{1}$ and $v_{2}$, one considers the set $E_{v_{1},v_{1}^{\prime}}$ which is the set of all $v_{2}^{\prime}$ which could be the velocity after collision of one particle, if the other velocity after collision is $v_{1}^{\prime}$ and one of the velocities before collision is $v_{1}$. See Fig. 3. The set $E_{v_{1},v_{1}^{\prime}}$ is the hyperplane orthogonal to $v_{1}-v_{1}^{\prime}$ and containing $v_{1}$ and we are writing it as the set
$E_{v_{1},v_{1}^{\prime}}:=\{v_{2}^{\prime}\in\mathds{R}^{3}\ :\ (v_{2}^{\prime}-v_{1})\cdot(v_{1}-v_{1}^{\prime})=0\},$ (5.1)
where $(v_{2}-v_{1})\cdot(v_{1}-v_{1}^{\prime})$ is the usual inner product such that
$(v_{2}-v_{1})\cdot(v_{1}-v_{1}^{\prime})=\langle v_{2}-v_{1},v_{1}-v_{1}^{\prime}\rangle=|v_{2}-v_{1}||v_{1}-v_{1}^{\prime}|\cos(v_{2}-v_{1},v_{1}-v_{1}^{\prime}).$
See [PW97, Section 3] for a further discussion of this.
$v_{1}^{\prime}$$v_{1}$$v_{2}$$v_{2}^{\prime}$$E_{v_{1},v_{1}^{\prime}}$$\omega$$\mathbb{S}_{v_{1},v_{2}}$
Figure 3. Picture of the hyperplane $E_{v_{1},v_{1}^{\prime}}$.
In order to represent the gain term with the Carleman representation, we set $v_{1}^{\prime}=v_{1}+q\omega$, $q\in\mathds{R}$ and $\omega\in\mathbb{S}^{2}$. Now, for any fixed $\omega$ and $v_{1}$ we can write $v_{2}^{\prime}=v_{2}-q\omega$, so $v_{2}=v_{2}^{\prime}+q\omega$. Then, with the change of variables $v_{2}=v_{2}^{\prime}+q\omega$, $\mathrm{d}v_{2}=\mathrm{d}v_{2}^{\prime}\mathrm{d}q$. Furthermore, by taking the polar coordinates of $v_{1}^{\prime}$ centred at $v_{1}$ we get
@@
$Q^{+}(f)(v_{1})=2\int_{\mathds{R}^{3}}\frac{f(v_{1}^{\prime})}{|v_{1}-v_{1}^{\prime}|^{2}}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})\ \mathcal{B}(v_{1}-v_{2},\omega)\,\mathrm{d}v_{2}^{\prime}\,\mathrm{d}v_{1}^{\prime}.$
For an extended explanation of the change of variables see [Silv16, Appendix A]. Now in order to have the variables $v_{1}^{\prime}$, $v_{2}^{\prime}$ inside the collision kernel $\mathcal{B}$, we use the identity $v_{1}-v_{2}=2v_{1}-v_{1}^{\prime}-v_{2}^{\prime}$ which comes from the conservation of momentum and we take
$Q^{+}(f)(v_{1})=2\int_{\mathds{R}^{3}}\frac{f(v_{1}^{\prime})}{|v_{1}-v_{1}^{\prime}|^{2}}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})\ \mathcal{B}(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}\,\mathrm{d}v_{1}^{\prime}.$
@@
$Q(f)(v_{1})$ $:=\int_{\mathds{R}^{3}}[\sigma(v_{1},v_{1}^{\prime})f(v_{1}^{\prime})-\sigma(v_{1}^{\prime},v_{1})f(v_{1})]\,\mathrm{d}v_{1}^{\prime}=K(f)(v_{1})-\nu(v_{1})f(v_{1}),$ (5.2)
@@
$=\int_{\mathds{R}^{3}}\left[\frac{2}{|v_{1}-v_{1}^{\prime}|^{2}}\int_{E_{v_{1},v_{1}^{\prime}}}g_{0}(v_{2}^{\prime})\mathcal{B}(2v_{1}-v_{1}^{\prime}-v_{2}^{\prime},\omega)\,\mathrm{d}v_{2}^{\prime}\right]f(v_{1}^{\prime})\,\mathrm{d}v_{1}^{\prime}$ (5.3)
@@
5.1. Estimate of the loss term of the collision operator
In this subsection, we analyse the behaviour of the loss term $\nu$ of the Boltzmann collision operator $Q$ as the velocity magnitude $|v_{1}|\to\infty$. Our goal is to establish Proposition 5.2, which characterizes the asymptotic decay of $\nu$ in the high-velocity regime. To proceed, we first state the following lemma which uses the notion of unimodal distribution. Unimodality refers to a distribution that has a maximum at 0 and is decaying in any radial direction.
Lemma 5.1.